News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 21 July 2000 4.00pm EDT

Several major solar flares and coronal mass ejections between July 12 and 14 caused enhanced particle rates in the Earth's vicinity lasting through July 16. We stopped execution of the on-board observing program at about 5pm on July 14, while all instruments were in a safe configuration for radiation belt passage. All observations on July 15 and 16 were postponed, and will be rescheduled as soon as possible. Following several team telecons through this time period, we determined that it was safe to resume observations upon exit from the radiation belts on Jul 17, and Chandra completed the observing schedule from that time on as planned. There is no apparent degradation of any science instrument or spacecraft system due to the solar particles.

A CAP was run Saturday night to unload momentum, which would have been done by the on-board stored commands if the observing program had not been stopped.

Processing of data showed that the observations of COMET c/LINEAR 1999 were successful: Chandra pointed correctly so as to catch this moving object, and the comet can be seen in the series of eight ACIS images which were obtained.

A routine real-time procedure was carried out to dump the CPE warm start counter, which showed the correct value of (1), as expected.

The HETG was successfully inserted via stored commands, the first time this had been attempted since failure of the B-side retract limit switch on June 13.

The planned observations for the coming week are as follows.

------------------------------------------
TERZAN 5                ACIS-I          Jul 24
SGR 1806-20             ACIS-S
1746-331                ACIS-S
Radiation Belt
G347.5-0.5              ACIS-I          Jul 25
G347.5-0.5              ACIS-I
4U 1543-47              ACIS-I          Jul 26
ABELL 3266              ACIS-I
CL J0542.8-4100         ACIS-I
Radiation Belt                          Jul 27
A2597                   ACIS-S          Jul 28
ABELL 496               ACIS-S
Q0054+144               ACIS-S
M31 TRANSIENT           ACIS-I
A2204                   ACIS-S
KES 73                  ACIS-S          Jul29
TERZAN 5                ACIS-I
A2163                   ACIS-I
ABELL 1995              ACIS-S
M 87                    ACIS-S
Radiation Belt                          Jul 30
M 87                    ACIS-S
1WGA1226.9+3332         ACIS-S

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports